If the pOH of solution is 1.7 then the [H+] of the solution will be what?
1
Expert's answer
2019-05-13T04:59:57-0400

"pH=14-pOH=14-1.7=12.3"

"[H+]=10^{-pH}=10^{-12.3}=5.012*10^{-13}"